OpenStudy (anonymous):

2^301 divide by 5, what is the least positive reminder?

Parth (parthkohli):

Try pattern recognition. \(2^1\) divided by 5: 2 \(2^2\): 4 \(2^3\): 3 \(2^4\): 1 \(2^5\): 2 . . .

Parth (parthkohli):

So basically, the remainder is 1 for all powers divisible by 4.

Parth (parthkohli):

300 is divisible by 4, so it would leave a remainder of 1. The series of remainders goes like this: 1, 2, 4, 3, 1, 2, 4, 3. So 2 always comes after 1.

Parth (parthkohli):

Do you understand?

OpenStudy (anonymous):

nope :(

Parth (parthkohli):

OK, let me explain clearly this time. I was a little messy the last time.

Parth (parthkohli):

\(2^0\) divided by \(5\) leaves remainder: 1. \(2^1\) divided by \(5\) leaves remainder: 2. \(2^2\) divided by \(5\) leaves remainder: 4. \(2^3\) divided by \(5\) leaves remainder: 3. \(2^4\) divided by \(5\) leaves remainder: 1. \(2^5\) divided by \(5\) leaves remainder: 2. \(2^6\) divided by \(5\) leaves remainder: 4. \(2^7\) divided by \(5\) leaves remainder: 3. \(2^8\) divided by \(5\) leaves remainder: 1.

Parth (parthkohli):

Do you see how the cycle is repeating? The remainders are coming out to be 1, 2, 4, 3, 1, 2, 4, 3, 1... Also, the remainder comes out to be 1 when the power of 2 is divisible by 4.

Parth (parthkohli):

So we have come up with a pattern and we can safely say that when the power of 2 is divisible by 4, the remainder is 1! Now 300 is divisible by 4, right? So what would you get as the remainder when you divide \(\large 2^{300}\) by 5?

OpenStudy (anonymous):

1

Parth (parthkohli):

Very good. Now, do you see how the power that comes after the powers divisible by 4 left a remainder of 2?

OpenStudy (anonymous):

ya

Parth (parthkohli):

Look at the pattern. \(2^1\) came right after \(2^0\) and left the remainder of 2. \(2^5\) came right after \(2^4\) and left the remainder of 2.

Parth (parthkohli):

So what can you say about \(2^{301}\) which comes right after \(2^{300}\)?

OpenStudy (anonymous):

reminder is 2

Parth (parthkohli):

Exactly. Also, if I may ask, what class are you in?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

can u answer this on the basis of binomial theorem?

OpenStudy (anonymous):

can u answer this on the basis of binomial theorem? @ganeshie8

ganeshie8 (ganeshie8):

try to write it as (x-1)^n or (x+1)^n and expand

OpenStudy (anonymous):

2(65-1)^50

ganeshie8 (ganeshie8):

expand

OpenStudy (anonymous):

2M(5)-2

OpenStudy (anonymous):

\[ 2^{301} = (5-3)^{301} = \sum_{k=0}^{301}5^k(-3)^{n-k} \]The only term which will not be divisible by \(5\) is the one where \(k=0\).

ganeshie8 (ganeshie8):

\(\large (x+y)^n = \sum \limits_{k=0}^n \binom{n}{k} x^{n-k}y^k\)

ganeshie8 (ganeshie8):

wio has it !

OpenStudy (anonymous):

so what is the reminder?

OpenStudy (anonymous):

Well, when \(k=0\) you get \[ (-3)^{301} \]

OpenStudy (anonymous):

so?

ganeshie8 (ganeshie8):

2(65-1)^50 after expanding, you should get : 2[M(5) + 1] right ?

OpenStudy (anonymous):